If the sum of two numbers is 30 and one of the nubers is x, write the product of the two numbers as a function of x.
This is what I had: n+x=30
x=30-n
f(x)=(30-n)(n)

Answer by scott8148(6628)

you're on the right track, but you want a function of x not n ... just swap the variables

n+x=30 ... n=30-x ... f(x)=xn ... f(x)=x(30-x)